Olive Company makes silver belt buckles. The company’s master budget appears in the first column of the table.

Required:
Complete the table by preparing Olive’s flexible budget for 5,200, 7,200, and 8,200 units. (Round your intermediate calculations to 2 decimal places.)

Master Budget (6,200 Units) Flexible Budget (5,200 Units) Flexible Budget (7,200 Units) Flexible Budget (8,200 Units)
Direct materials $1,860
Direct labor 2,480
Variable manufacturing overhead 2,480
Fixed manufacturing overhead 17,600
Total manufacturing cost $24,420
Answer #1

NOTE:Variable cost per unit and Total fixed cost do not change with change in units

Direct material/unit=(1860/6200)=$0.3

Direct labor=(2480/6200)=0.4

Variable manufacturing overhead=(2480/6200)=0.4

Master Budget (6,200 Units) Flexible Budget (5,200 Units) Flexible Budget (7,200 Units) Flexible Budget (8,200 Units)
Direct materials $1,860 1560 2160 2460
Direct labor 2,480 2080 2880 3280
Variable manufacturing overhead 2,480 2080 2880 3280
Fixed manufacturing overhead 17,600 17600 17600 17600
Total manufacturing cost $24,420 23320 25520 26620